# Superconductivity in Electron Liquids: Precision Many-Body Treatment of Coulomb Interaction

### v1

@@ II.3 Precursory Cooper Flow
The anomalous vertex function, $\Lambda_{\mathbf{k}\omega}$, is used not only to predict $T_{\text{c}}$ but also to study pairing fluctuations in the normal state and the critical gap function $\Delta_{\mathbf{k}\omega}$. We denote the low-frequency limit of $\Lambda_{\mathbf{k}\omega}$ averaged over the Fermi surface as $\Lambda_{0}$. It obeys a universal scaling relation known as the precursory Cooper flow (PCF) [pengcheng]:
@@
where both the dimensionless coupling constant $g$ and high-energy cutoff $\omega_{\Lambda}$ depend on the microscopic details of the system. For negative values of $g$, the vertex function diverges at $T_{\text{c}}=\omega_{\Lambda}e^{1/g}$, signaling the onset of Cooper instability. By computing $\Lambda_{0}$ at several low-temperature points above $T_{\text{c}}$ and extrapolating the data according to the PCF scaling, one can accurately predict $T_{\text{c}}$ without the need to perform computationally challenging calculations at $T_{c}$ or in the symmetry-broken phase below $T_{\text{c}}$.
@@
is the precise scaling law (10). In a system with strong repulsive interactions, there is no simple way to reliably extrapolate $h(T)$ towards low temperature [implicit].
@@ III.1 Theoretical Overview
Eliashberg [eliashberg1960interactions, eliashberg1961temperature] recognized that a discussion of superconductivity in conventional materials should focus on interactions near the Fermi surface and proposed a “downfolding” approximation based on effective interactions between low-energy electronic states that led to a more tractable frequency-only equation for the superconducting gap function. However, as discussed extensively above, to date the parameters (especially those relating to the electron-electron interactions) in the downfolded theory have typically been treated in ways that are phenomenological or theoretically unjustified.
The Wilsonian renormalization scheme is a formally rigorous method for downfolding an interacting electron problem to obtain a low energy theory [RAINER1986371, Shankar94]. It formalizes Landau Fermi liquid theory by partitioning electron and phonon propagators into high-energy and low-energy contributions separated by an intermediate energy scale $\omega_{\text{D}}\ll\omega_{\text{c}}\ll E_{F}$. The high-energy contributions are then absorbed into effective couplings that govern the low-energy theory for the quasiparticle pair-field. However, a crucial complication arises in the standard applications of this scheme to the electron liquid: the low-energy electrons, which are essential for dynamic screening of the long-range Coulomb interaction, are absent in the renormalization of the effective coupling because the separation is constructed in the particle-hole channel. As a result, the Coulomb interaction remains unscreened below the energy scale $\omega_{c}$, leading to a singularity in the effective coupling that is awkward to handle.
In this work, we introduce the energy scale separation in the two-electron channel rather than in the single-electron or electron-hole channels. This approach preserves key features of the screening while avoiding the subtle issues encountered in previous attempts [RAINER1986371, Shankar94]. We decompose the pair propagator into low-energy (IR) coherent and high-energy (UV) incoherent components, separated at the energy scale $\omega_{\text{c}}$,
@@
Here, $\mathbf{k},\omega$ is the momentum/frequency of the pair-field, $\phi$ denotes the incoherent contribution, an intrinsic property of the electron liquid with phonon contributions $\mathcal{O}\left(\frac{\omega_{D}}{E_{F}}\right)$ which remains regular across the Fermi surface at low temperature. We have anticipated subsequent results by writing the coherent contribution in terms of $z^{\text{e}}$, the quasiparticle weight arising from e-e interactions which is independent of frequency and momentum on the scales of interest, and $z_{\omega}^{\text{ph}}$, the frequency-dependent quasiparticle weight from e-ph interaction. $\epsilon_{\mathbf{k}}$ is the linearized quasiparticle dispersion.
@@
where the notation $\mu^{*}=\mu_{\omega_{\text{c}}},\,\omega_{*}=\omega_{\text{c}}$ is used to present it in the familiar ME form.
@@ III.2 The pseudopotential
Our results for $\mu_{{}_{E_{\text{F}}}}$ in UEG as a function of $r_{\text{s}}$ are presented in Fig. 4; the corresponding numerical values are listed in Table 1. We find $\mu_{{}_{E_{\text{F}}}}$ to be significantly larger (by a factor of three at $r_{\text{s}}=5$) compared to MA and static RPA estimates [morel_anderson, RPA_mu, RPA_mu2]. Even more profound is the disagreement with dynamic RPA that predicts negative $\mu_{\omega_{\text{c}}}$ values for $r_{\text{s}}>2$ [takada1978plasmon, takada1993, rietschel, xiansheng, 23tao]. This outcome clearly demonstrates that the MA and RPA approximations are out of control at $r_{\text{s}}>0.5$, which one could have expected from the mere fact that at $r_{\text{s}}>0.5$ the results of static and dynamic RPA dramatically differ from each other.
Figure 4: Dimensionless “bare” Coulomb pseudopotential, $\mu_{E_{F}}$, as a function of $r_{\text{s}}$ for the 3D UEG extracted from VDiagMC data for $\mu_{\omega_{\text{c}}}$ by inverting relation (25); the solid line represents a linear fit to the VDiagMC data. The exact VDiagMC values corresponding to these integer $r_{\text{s}}$ points are listed in Table 1. At $r_{\text{s}}>0.5$, the VDiagMC results demonstrate a dramatic deviation from predictions of three standard approximations: static random phase approximation (RPA) $\mu_{\text{RPA-static}}$ (red dashed curve), $\mu_{\text{MA}}$ (green dashed curve) based on the Yukawa interaction with Thomas-Fermi screening momentum, and the dynamic RPA (see, e.g., Ref. [pengcheng]), $\mu_{\text{RPA-dynamic}}$. Note perfect agreement between all curves at $r_{\text{s}}\ll 1$.
@@
In addition to parameterizing the value of the Coulomb pseudopotential as a function of low frequency cutoff, Eq. (25), the bare pseudopotential $\mu_{{}_{E_{\text{F}}}}$ plays yet another role that is closely related to superconductivity. The low-temperature response of the normal-Fermi-liquid state to the uniform pair-creating perturbation, $\chi_{0}$, has the form (see, e.g., Ref. [pengcheng]) $\chi_{0}\,\propto\,{z_{\rm e}^{2}\over m^{*}\mu_{{}_{E_{\text{F}}}}}$ for $\mu_{{}_{E_{\text{F}}}}>0,\;\omega_{D}\ll T\ll E_{\text{F}},\omega_{p}).$ From our results for $\mu_{{}_{E_{\text{F}}}}$ we thus conclude that $\chi_{0}$ gets substantially suppressed with increasing $r_{\rm s}$.
@@ III.3 Validation of the Fermi-Surface Downfolding
To validate this approximation for the systems under consideration, we investigate a toy model with the two-particle-irreducible vertex function, $\tilde{\Gamma}^{\text{e}}$, approximated by the dynamically screened Coulomb interaction of the uniform electron gas within the RPA, $W_{\text{RPA}}=v_{q}/(1-v_{q}\Pi^{0}_{q\nu})$. For the effective phonon-mediated interaction, $W^{\text{ph}}_{q\nu}$, we consider a model form from Ref. [23tao]:
@@
While the downfolding framework is clearly robust in this physical regime, its accuracy could be challenged in extreme cases where the separation of energy scales is less pronounced. For instance, downfolding can become questionable in systems with low electron densities where $E_{\text{F}}\sim\omega_{\text{D}}$ (a regime outside the scope of this discussion), as well as in dense systems with small $r_{\text{s}}$. This includes certain astrophysical objects, such as the interiors of white dwarfs (slowly evolving into black dwarfs). In these systems, electrons may become superconducting at sufficiently low temperatures [ginzburg_superconductivity_1968, trubnikov1968white], and their $r_{\text{s}}$ values can be as small as 0.01. This leads to soft plasmon frequencies and potential problems with the accuracy of the downfolding approximation. Similar, and potentially more severe, problems are anticipated in two-dimensional electron gas (2DEG) systems [katsnelson2023screening], provided the Coulomb interaction is not screened by a substrate. In that case, the plasmon mode remains gapless across all density regimes, a situation that warrants further careful investigation. A detailed analysis of such exotic regimes goes beyond the scope of the present work.
@@ IV.1 Uniform Electron Gas
Despite the relative simplicity of the UEG model (compared to real materials) precise calculation of its Coulomb pseudopotential presents a significant challenge. We are not aware of any calculation performed with controlled accuracy for $r_{\text{s}}>1$. The most challenging aspect is the evaluation of the four-point vertex function in the Cooper channel on the Fermi surface. In the strongly correlated regime, this quantity cannot be obtained by traditional ground-state methods such as variational and diffusion Monte Carlo [vmc1, vmc2, dmc1].
In this context, the recently developed method of Variational diagrammatic Monte Carlo [diagmc1, diagmc2, diagmc2010, detdiagmc, diagmclatest, diagmckozik, vdiagmc1, vdiagmc2] (VDiagMC) emerges as the right tool to address the problem. Unlike the majority of quantum Monte Carlo techniques, VDiagMC is based on stochastic sampling of Feynman diagrams to high order and provides direct access to the vertex function needed for evaluation of $\mu^{*}$. The key principle of VDiagMC is optimization of the starting point to set up the diagrammatic expansion order-by-order. As a result, one obtains accurate converged answers even beyond the weak-coupling regime.
@@
As detailed in Ref. [hou_feynman_2024], the efficiency of VDiagMC is greatly enhanced by representing high-order Feynman diagrams of vertex functions as computational graphs. This approach leverages the structure of Dyson-Schwinger and Parquet equations to express diagrams in a compressed form with a fractal structure of tensor operations, significantly reducing computational redundancy. Notably, this computational graph representation allows for highly efficient implementation of field-theoretic renormalization schemes of the bare parameters using Taylor-mode automatic differentiation algorithms [Taylor1, Taylor2, Taylor3], reducing the computational cost of evaluating the renormalized set of diagrams from exponential to sub-exponential scaling.
Building upon these techniques, we developed a Feynman diagram compiler that generates, optimizes, and converts the renormalized diagrammatic series into a compressed computational graph across various platforms using machine learning frameworks. This enables the implementation of high-dimensional Monte Carlo integration algorithms with state-of-the-art importance sampling techniques, such as the VEGAS [vegas, vegasenhanced] adaptive algorithm and more advanced methods employing normalizing flow neural networks [normalization_flow]. Our codes support both the conventional CPU clusters and GPU platforms, allowing for scalable and efficient computations. In practice, we can reach sufficiently high expansion orders to accurately determine various quasiparticle properties, including density, spin susceptibility, and effective mass. This combination of computational graph representation, Taylor-mode automatic differentiation, and modern machine learning frameworks provides a robust and efficient AI Tech Stack for quantum field theory.
@@ IV.2 Coulomb Pseudopotential: Homotopic expansion
At low temperature, the convergence of the series at the physical value $\xi=1$ becomes problematic due to the $(\ln T)^{N}$ scaling of the $N$-th order term arising from nested particle-particle bubbles in the diagrammatic expansion. The solution to the convergence problem comes with a homotopic trick (cf. Ref. [homotopic_expansions]). Assuming analyticity of $\gamma_{T}(\xi)$ as a function of $\xi$, we want to construct a matching temperature-independent analytic function $\mu_{\omega_{c}}(\xi)$ such that, on the one hand, it could be analytically expressed in terms of $\gamma_{T}(\xi)$ and, on the other hand, would feature the homotopic requirement that $\mu_{\omega_{c}}(\xi=1)$ be equal to the value of $\mu_{\omega_{c}}$ defined by Eq. (23). The series (28) will then be readily converted into the homotopic expansion
@@ V Electron-Phonon Coupling from Band Theory
Accurate determination of the phonon-mediated electron-electron attraction quantified by the dimensionless coupling constant $\lambda$ is fundamental to understanding and predicting phonon-mediated SC. Current theories of conventional superconductivity are based on defining electron-phonon coupling by using density functional theory to estimate the response of the electronic states to changes in atomic positions. This response is efficiently implemented in the density functional perturbation theory (DFPT) computational method [dfpt_rev, EPW1, EPW2, giustino_2017]. Currently, DFPT underlies most ab initio predictions of conventional superconductivity, yet its accuracy in correlated systems remains untested. Moreover, precise benchmarks help establish semi-phenomenological pathways for systematically correcting DFPT results, potentially extending its applicability to strongly correlated superconductors. In this section we address the fundamental question: under what conditions and with what accuracy does the EFT framework agree with the results of DFPT [dfpt_rev, EPW1, EPW2, giustino_2017] ?
@@
$\lambda=N_{\text{F}}\sum_{\kappa}\left\langle\frac{g_{\kappa}^{2}(\mathbf{k},\mathbf{q})}{\omega^{2}_{\kappa,\mathbf{q}}}\right\rangle_{FS}\,.$ (31)
@@
To obtain $\lambda$, the effective e-ph coupling must be combined with the physical phonon spectrum and density of states, $N_{\text{F}}$. This is where the two methods are radically different: EFT is based on the quasiparticle density of states, while DFPT uses the band-structure density of states (bare one for UEG), $N^{(0)}_{\text{F}}$. The two differ by the ratio of the quasiparticle mass to the electron mass. Recent high-precision calculations for UEG [vdiagmc2, markus_2023] find this ratio to be very close to unity (with sub-percent accuracy at $r_{\text{s}}=5$). Therefore, the two densities of states end up nearly identical despite strong correlations.

### v2

@@ II.3 Precursory Cooper Flow
The anomalous vertex function, $\Lambda_{\mathbf{k}\omega}$, is used not only to predict $T_{\text{c}}$ but also to study pairing fluctuations in the normal state and the critical gap function $\Delta_{\mathbf{k}\omega}$. We denote the low-frequency limit of $\Lambda_{\mathbf{k}\omega}$ averaged over the Fermi surface as $\Lambda_{0}$. It obeys a universal scaling relation known as the precursory Cooper flow (PCF) [44]:
@@
where both the dimensionless coupling constant $g$ and high-energy cutoff $\omega_{\Lambda}$ depend on the microscopic details of the system. For negative values of $g$, the vertex function diverges at $T_{\text{c}}=\omega_{\Lambda}e^{1/g}$, signaling the onset of Cooper instability. By computing $\Lambda_{0}$ at several low-temperature points above $T_{\text{c}}$ and extrapolating the data according to the PCF scaling, one can accurately predict $T_{\text{c}}$ without the need to perform computationally challenging calculations at $T_{\text{c}}$. This is particularly advantageous when complex frequency-dependent interactions prevent standard linearized ME solutions from extrapolating reliably from higher temperatures [19, 44], which would otherwise force one to solve the equations directly at the critical point.
@@
is the precise scaling law (10). In a system with strong repulsive interactions, there is no simple way to reliably extrapolate $h(T)$ towards low temperature [19].
@@ III.1 Theoretical Overview
Eliashberg [22, 23] recognized that a discussion of superconductivity in conventional materials should focus on interactions near the Fermi surface and proposed a “downfolding” approximation based on effective interactions between low-energy electronic states that led to a more tractable frequency-only equation for the superconducting gap function. However, as discussed extensively above, to date the parameters (especially those relating to the electron-electron interactions) in the downfolded theory have typically been treated in ways that are phenomenological or theoretically unjustified.
The Wilsonian renormalization scheme is a formally rigorous method for downfolding an interacting electron problem to obtain a low energy theory [83, 95]. It formalizes Landau Fermi liquid theory by partitioning electron and phonon propagators into high-energy and low-energy contributions separated by an intermediate energy scale $\omega_{\text{D}}\ll\omega_{\text{c}}\ll E_{F}$. The high-energy contributions are then absorbed into effective couplings that govern the low-energy theory for the quasiparticle pair-field. However, a crucial complication arises in the standard applications of this scheme to the electron liquid: the low-energy electrons, which are essential for dynamic screening of the long-range Coulomb interaction, are absent in the renormalization of the effective coupling because the separation is constructed in the particle-hole channel. As a result, the Coulomb interaction remains unscreened below the energy scale $\omega_{c}$, leading to a singularity in the effective coupling that is awkward to handle.
In this work, we introduce the energy scale separation in the two-electron channel rather than in the single-electron or electron-hole channels. This approach preserves key features of the screening while avoiding the subtle issues encountered in previous attempts [83, 95]. We decompose the pair propagator into low-energy (IR) coherent and high-energy (UV) incoherent components, separated at the energy scale $\omega_{\text{c}}$,
@@
Here, $\mathbf{k},\omega$ is the momentum/frequency of the pair-field, $\phi$ denotes the incoherent contribution, an intrinsic property of the electron liquid with phonon contributions $\mathcal{O}\left(\frac{\omega_{D}}{E_{F}}\right)$ which remains regular across the Fermi surface at low temperature. We have anticipated subsequent results by writing the coherent contribution in terms of $z^{\text{e}}$, the quasiparticle weight arising from e-e interactions which is independent of frequency and momentum on the scales of interest, and $z_{\omega}^{\text{ph}}$, the frequency-dependent quasiparticle weight from e-ph interaction. $\varepsilon_{\mathbf{k}}$ is the linearized quasiparticle dispersion renormalized by e-e interactions. Formally, the effective mass renormalization $m^{*}$ is fully incorporated into $\varepsilon_{\mathbf{k}}$ and the subsequent effective density of states $N_{F}^{*}$. Although established QMC and DiagMC results [38, 42] indicate that $m^{*}$ deviates only slightly from the bare mass in the density range of interest, we retain this distinction to separate mass renormalization from the spectral weight reduction $z^{\text{e}}$.
@@
where the notation $\mu^{*}=\mu_{\omega_{\text{c}}},\,\omega_{*}=\omega_{\text{c}}$ is used to present it in the familiar ME form. We note that in the low-frequency limit, the term $\lambda_{\omega\omega^{\prime}}z^{\text{ph}}_{\omega^{\prime}}$ reduces to $\lambda/(1+\lambda)$.
@@ III.2 The pseudopotential
Our results for $\mu_{{}_{E_{\text{F}}}}$ in UEG as a function of $r_{\text{s}}$ are presented in Fig. 4; the corresponding numerical values are listed in Table 1. We find $\mu_{{}_{E_{\text{F}}}}$ to be significantly larger (by a factor of three at $r_{\text{s}}=5$) compared to MA and static RPA estimates [69, 91, 2]. Even more profound is the disagreement with dynamic RPA that predicts negative $\mu_{\omega_{\text{c}}}$ values for $r_{\text{s}}>2$ [100, 101, 89, 15, 108]. This outcome clearly demonstrates that the MA and RPA approximations are out of control at $r_{\text{s}}>0.5$, which one could have expected from the mere fact that at $r_{\text{s}}>0.5$ the results of static and dynamic RPA dramatically differ from each other.
Figure 4: Dimensionless “bare” Coulomb pseudopotential, $\mu_{E_{F}}$, as a function of $r_{\text{s}}$ for the 3D UEG extracted from VDiagMC data for $\mu_{\omega_{\text{c}}}$ by inverting relation (25); the solid line represents a linear fit to the VDiagMC data. The exact VDiagMC values corresponding to these integer $r_{\text{s}}$ points are listed in Table 1. At $r_{\text{s}}>0.5$, the VDiagMC results demonstrate a dramatic deviation from predictions of three standard approximations: static random phase approximation (RPA) $\mu_{\text{RPA-static}}$ (red dashed curve), $\mu_{\text{MA}}$ (green dashed curve) based on the Yukawa interaction with Thomas-Fermi screening momentum, and the dynamic RPA (see, e.g., Ref. [44]), $\mu_{\text{RPA-dynamic}}$. Note perfect agreement between all curves at $r_{\text{s}}\ll 1$.
@@
In addition to parameterizing the value of the Coulomb pseudopotential as a function of low frequency cutoff, Eq. (25), the bare pseudopotential $\mu_{{}_{E_{\text{F}}}}$ plays yet another role that is closely related to superconductivity. The low-temperature response of the normal-Fermi-liquid state to the uniform pair-creating perturbation, $\chi_{0}$, has the form (see, e.g., Ref. [44]) $\chi_{0}\,\propto\,{z_{\rm e}^{2}\over m^{*}\mu_{{}_{E_{\text{F}}}}}$ for $\mu_{{}_{E_{\text{F}}}}>0,\;\omega_{D}\ll T\ll E_{\text{F}},\omega_{p}).$ From our results for $\mu_{{}_{E_{\text{F}}}}$ we thus conclude that $\chi_{0}$ gets substantially suppressed with increasing $r_{\rm s}$.
@@ III.3 Validation of the Fermi-Surface Downfolding
To validate this approximation for the systems under consideration, we investigate a toy model with the two-particle-irreducible vertex function, $\tilde{\Gamma}^{\text{e}}$, approximated by the dynamically screened Coulomb interaction of the uniform electron gas within the RPA, $W_{\text{RPA}}=v_{q}/(1-v_{q}\Pi^{0}_{q\nu})$. For the effective phonon-mediated interaction, $W^{\text{ph}}_{q\nu}$, we consider a model form from Ref. [108]:
@@
While the downfolding framework is clearly robust in this physical regime, its accuracy could be challenged in extreme cases where the separation of energy scales is less pronounced. For instance, downfolding can become questionable in systems with low electron densities where $E_{\text{F}}\sim\omega_{\text{D}}$ (a regime outside the scope of this discussion), as well as in dense systems with small $r_{\text{s}}$. This includes certain astrophysical objects, such as the interiors of white dwarfs (slowly evolving into black dwarfs). In these systems, electrons may become superconducting at sufficiently low temperatures [31, 104], and their $r_{\text{s}}$ values can be as small as 0.01. This leads to soft plasmon frequencies and potential problems with the accuracy of the downfolding approximation. Similar, and potentially more severe, problems are anticipated in two-dimensional electron gas (2DEG) systems [49], provided the Coulomb interaction is not screened by a substrate. In that case, the plasmon mode remains gapless across all density regimes, a situation that warrants further careful investigation. A detailed analysis of such exotic regimes goes beyond the scope of the present work.
@@ IV.1 Uniform Electron Gas
Despite the relative simplicity of the UEG model (compared to real materials) precise calculation of its Coulomb pseudopotential presents a significant challenge. We are not aware of any calculation performed with controlled accuracy for $r_{\text{s}}>1$. The most challenging aspect is the evaluation of the four-point vertex function in the Cooper channel on the Fermi surface. In the strongly correlated regime, this quantity cannot be obtained by traditional ground-state methods such as variational and diffusion Monte Carlo [66, 17, 85].
In this context, the recently developed method of Variational diagrammatic Monte Carlo [80, 81, 107, 90, 106, 56, 18, 39] (VDiagMC) emerges as the right tool to address the problem. Unlike the majority of quantum Monte Carlo techniques, VDiagMC is based on stochastic sampling of Feynman diagrams to high order and provides direct access to the vertex function needed for evaluation of $\mu^{*}$. The key principle of VDiagMC is optimization of the starting point to set up the diagrammatic expansion order-by-order. As a result, one obtains accurate converged answers even beyond the weak-coupling regime.
@@
As detailed in Ref. [45], the efficiency of VDiagMC is greatly enhanced by representing high-order Feynman diagrams of vertex functions as computational graphs. This approach leverages the structure of Dyson-Schwinger and Parquet equations to express diagrams in a compressed form with a fractal structure of tensor operations, significantly reducing computational redundancy. Notably, this computational graph representation allows for highly efficient implementation of field-theoretic renormalization schemes of the bare parameters using Taylor-mode automatic differentiation algorithms [12, 35, 102], reducing the computational cost of evaluating the renormalized set of diagrams from exponential to sub-exponential scaling.
Building upon these techniques, we developed a Feynman diagram compiler that generates, optimizes, and converts the renormalized diagrammatic series into a compressed computational graph across various platforms using machine learning frameworks. This enables the implementation of high-dimensional Monte Carlo integration algorithms with state-of-the-art importance sampling techniques, such as the VEGAS [74, 57] adaptive algorithm and more advanced methods employing normalizing flow neural networks [14]. Our codes support both the conventional CPU clusters and GPU platforms, allowing for scalable and efficient computations. In practice, we can reach sufficiently high expansion orders to accurately determine various quasiparticle properties, including density, spin susceptibility, and effective mass. This combination of computational graph representation, Taylor-mode automatic differentiation, and modern machine learning frameworks provides a robust and efficient AI Tech Stack for quantum field theory.
@@ IV.2 Coulomb Pseudopotential: Homotopic expansion
At low temperature, the convergence of the series at the physical value $\xi=1$ becomes problematic due to the $(\ln T)^{N}$ scaling of the $N$-th order term arising from nested particle-particle bubbles in the diagrammatic expansion. The solution to the convergence problem comes with a homotopic trick (cf. Ref. [51]). Assuming analyticity of $\gamma_{T}(\xi)$ as a function of $\xi$, we want to construct a matching temperature-independent analytic function $\mu_{\omega_{c}}(\xi)$ such that, on the one hand, it could be analytically expressed in terms of $\gamma_{T}(\xi)$ and, on the other hand, would feature the homotopic requirement that $\mu_{\omega_{c}}(\xi=1)$ be equal to the value of $\mu_{\omega_{c}}$ defined by Eq. (23). The series (28) will then be readily converted into the homotopic expansion
@@ V Electron-Phonon Coupling from Band Theory
Accurate determination of the phonon-mediated electron-electron attraction quantified by the dimensionless coupling constant $\lambda$ is fundamental to understanding and predicting phonon-mediated SC. Current theories of conventional superconductivity are based on defining electron-phonon coupling by using density functional theory to estimate the response of the electronic states to changes in atomic positions. This response is efficiently implemented in the density functional perturbation theory (DFPT) computational method [9, 32, 78, 33]. Currently, DFPT underlies most ab initio predictions of conventional superconductivity, yet its accuracy in correlated systems remains untested. Moreover, precise benchmarks help establish semi-phenomenological pathways for systematically correcting DFPT results, potentially extending its applicability to strongly correlated superconductors. In this section we address the fundamental question: under what conditions and with what accuracy does the EFT framework agree with the results of DFPT [9, 32, 78, 33] ?
@@
$\lambda=N_{\text{F}}\sum_{\kappa}\left\langle\frac{g_{\kappa}^{2}(\mathbf{k},\mathbf{q})}{\omega^{2}_{\kappa,\mathbf{q}}}\right\rangle_{\text{FS}}\,.$ (31)
@@
To obtain $\lambda$, the effective e-ph coupling must be combined with the physical phonon spectrum and density of states, $N_{\text{F}}$. This is where the two methods are radically different: EFT is based on the quasiparticle density of states, while DFPT uses the band-structure density of states (bare one for UEG), $N^{(0)}_{\text{F}}$. The two differ by the ratio of the quasiparticle mass to the electron mass. Recent high-precision calculations for UEG [39, 43] find this ratio to be very close to unity (with sub-percent accuracy at $r_{\text{s}}=5$). Therefore, the two densities of states end up nearly identical despite strong correlations.